# 3DGS-LM: Faster Gaussian-Splatting Optimization with Levenberg-Marquardt

### v1

@@ 17      $\mathbf{x}_{i+1}{=}\mathbf{x}_{i}{+}\alpha_{i}\mathbf{p}_{i}$
18      $\mathbf{r}_{i+1}{=}\mathbf{r}_{i}{-}\alpha_{i}g_{i}$
@@ 21      $\mathbf{p}_{i+1}=\mathbf{z}_{i+1}+\beta_{i}\mathbf{p}_{i}$
Most of the work in every PCG iteration is consumed by calculating the matrix-vector product $\mathbf{g}_{i}{=}\mathbf{J}^{T}\mathbf{J}\mathbf{p}_{i}$. We implement this product in two stages by first calculating $\mathbf{u}_{i}{=}\mathbf{J}\mathbf{p}_{i}$ and then $\mathbf{g}_{i}{=}\mathbf{J}^{T}\mathbf{u}_{i}$. Calculating the non-zero values of $\mathbf{J}$ requires backpropagating from every residual through the $\alpha$-blending and splat projection steps back to all Gaussian parameters. The original tile-based rasterizer of 3DGS [20] implements the calculation of $\mathbf{J}^{T}\mathbf{x}$ with a per-pixel parallelization. That is, every thread handles one ray, stepping backwards along all splats that this ray hit. We found that this parallelization is too slow for an efficient PCG implementation. The reason is that we have to repeat this calculation multiple times: per PCG iteration we do it once for $\mathbf{u}_{i}$ and once for $\mathbf{g}_{i}$. As a consequence, we re-calculate the same intermediate $\alpha$-blending states, since the loop over splats is evaluated multiple times.
@@
Figure 3: Parallelization Strategy And Caching Scheme. We implement the PCG algorithm with efficient CUDA kernels, that use a gradient cache to calculate Jacobian-vector products. Before PCG starts, we create the gradient cache following the per-pixel parallelization of 3DGS [20]. Afterwards, we sort the cache by Gaussians to ensure coalesced read accesses. The cache decouples splats along rays, which allows us to parallelize per-pixel-per-splat during PCG.
Our key idea is to change the parallelization from per-pixel to per-pixel-per-splat. We summarize this pattern in Fig. 3. We describe the backward pass from a residual $r$ to a Gaussian parameter $x_{i}$ as:
$\frac{\partial r}{\partial x_{i}}=\frac{\partial r}{\partial p}\frac{\partial p}{\partial s}\frac{\partial s}{\partial x_{i}}$ (6)
@@
where $\frac{\partial r}{\partial p}$ is the gradient from the residual to the pixel, $\frac{\partial p}{\partial s}$ from the pixel to the projected splat, and $\frac{\partial s}{\partial x_{i}}$ from the splat to the Gaussian parameter. Calculating $\frac{\partial p}{\partial s}$, requires access to the accumulated transmittance and color in the $\alpha$-blending process for that pixel until splat $s$. Instead of looping over all splats along a ray and recalculating $\frac{\partial p}{\partial s}$ every time, we cache this gradient once (Fig. 3 left). Every time we need to calculate $\mathbf{u}_{i}$ or $\mathbf{g}_{i}$ in the PCG algorithm, we then read the corresponding gradients from the cache (Fig. 3 right). This allows us to directly parallelize over all splats in all pixels, which drastically accelerates the runtime since we no longer have to loop over rays.
Storing these gradients in a cache consumes additional GPU memory and is largely controlled by how many images (rays) we process in each PCG iteration and how many splats contribute to the final color along each ray. We propose an efficient subsampling scheme in Sec. 3.3.2 to limit the memory to the available budget.
3.3.1 CUDA Kernel Design
We describe the parallelization pattern and output of every CUDA kernel that we use to implement Algorithm 1. We refer to the supplemental material for more details.
buildCache We use the per-pixel parallelization to calculate all gradients $\frac{\partial p}{\partial s}$. For coalesced write accesses, we store the cache values sorted by pixels (Fig. 3 left). Additionally, we calculate $\mathbf{b}=-\mathbf{J}^{T}\mathbf{F}$ following the implementation of the differentiable rasterizer in 3DGS [20], by splitting this calculation into three smaller kernels, where we write out the cache completely in the first kernel.
sortCacheByGaussians All remaining kernels require the cache to be sorted over Gaussians to ensure coalesced read accesses. We sort the cache by Gaussians and calculate a prefix sum over the number of entries per Gaussian. Subsequent kernels use this to index into the cache.
diagJTJ We use the Jacobi preconditioner $\mathbf{M}^{-1}{=}1/\text{diag}(\mathbf{J}^{T}\mathbf{J})$ when running PCG (Algorithm 1). This kernel computes the required entries by utilizing the per-pixel-per-splat parallelization and the cache. Every thread directly squares its calculated gradients and atomically adds them to the output vector.
applyJ We compute $\mathbf{u}_{i}{=}\mathbf{J}\mathbf{p}_{i}$ by utilizing the per-pixel-per-splat parallelization and the cache. Additionally, we resort the input vector $\mathbf{p}_{i}$ for a coalesced memory access using the sortX kernel. Every thread sums up the calculated gradients and atomically adds them to the output vector.
applyJT We compute $\mathbf{g}_{i}{=}\mathbf{J}^{T}\mathbf{u}_{i}$ by utilizing the per-pixel-per-splat parallelization and the cache. Similar to buildCache, we split the computation in three kernels.
3.3.2 Image Subsampling Scheme
@@
Our caching data structure consumes additional GPU memory. For high resolution images in a dense reconstruction setup, the number of rays and thus the cache size can grow too large. To this end, we split the images into batches and solve the normal equations independently, following  Eq. 3. This allows us to store the cache only for one batch at a time. Concretely, for $n$ batches, we obtain $n$ update vectors and combine them in a weighted mean:
$\Delta=\sum_{i=1}^{n}\frac{\mathbf{M}_{i}\Delta_{i}}{\sum_{k=1}^{n}\mathbf{M}_{k}}$ (7)
@@
where we use the inverse of the PCG preconditioner $M_{i}{=}\text{diag}(\mathbf{J}_{i}^{T}\mathbf{J}_{i})$ as the weights. We refer to the supplement for a derivation of the weights. These weights balance the importance of update vectors across batches based on how much each Gaussian parameter contributed to the final colors in the respective images. This subsampling scheme allows us to control the cache size relative to the number of images in a batch. In practice, we can choose batch sizes between 25 and 70 images and use up to $n{=}4$ batches per LM iteration. We either select the images at random or, if the scene was captured along a smooth trajectory, in a strided fashion to maximize the scene coverage in every batch.
@@
Figure 4: Comparison of initialization iterations. In our first stage, we initialize the Gaussians with gradient descent for $K$ iterations, before finetuning with our LM optimizer. After 6000 or 8000 iterations, our method converges faster than the baseline. With less iterations, pure LM is slower, which highlights the importance of our two stage approach. Results reported on the “garden” scene from MipNeRF360 [28] without densification.
We utilize our LM implementation in the second stage of 3DGS optimization (see Fig. 2). Before that, we use the ADAM optimizer to obtain an initialization of the Gaussian parameters. It is also possible to use the LM optimizer from the beginning, however this does not bring any additional speed-up (see Fig. 4). In the beginning of optimization, gradient descent makes rapid progress by optimizing the Gaussians from a single image per iteration. In contrast, we sample many images in every LM iteration, which makes every iteration more time-consuming. This additional compute overhead is especially helpful to converge to optimal Gaussian parameters quicker (see Fig. 1 left).
Splitting the method in two stages also allows us to complete the densification of the Gaussians before employing the LM optimizer, which simplifies the implementation.
@@ 4 Results
Baselines We compare our LM optimizer against ADAM in multiple reference implementations of 3DGS. This shows, that our method is compatible with other runtime improvements. In other words, we can swap out the optimizer and retain everything else. Concretely, we compare against the original 3DGS implementation [20], its reimplementation “gsplat” [41], and DISTWAR [11]. Additionally, we compare against Taming-3DGS [27] by utilizing their released source code, that contains the rasterizer improvements and uses the 3DGS [20] densification heuristics (referred to as “Taming-3DGS” in the following). We run all baselines for 30K iterations with their default hyperparameters.
Datasets and Metrics We benchmark our runtime improvements on three established datasets: Tanks&Temples [24], Deep Blending [18], and MipNeRF360 [4]. These datasets contain in total 13 scenes that cover bounded indoor and unbounded outdoor environments. We fit all scenes for every method on the same NVIDIA A100 GPU using the train/test split as proposed in the original 3DGS [20] publication. To measure the quality of the reconstruction, we report peak signal-to-noise ratio (PSNR), structural similarity (SSIM), and perceptual similarity (LPIPS) [44] averaged over all test images. Additionally, we report the optimization runtime and the maximum amount of consumed GPU memory.
Implementation Details For our main results, we run the first stage for 20K iterations with the default hyperparameters of the respective baseline. The densification is completed after 15K iterations. Afterwards, we only have to run 5 LM iterations with 8 PCG iterations each to converge on all scenes. This showcases the efficiency of our optimizer. Since the image resolutions are different for every dataset, we select the batch-size and number of batches such that the consumed memory for caching is similar. We select 25 images in 4 batches for MipNeRF360 [4], 25 images in 3 batches for Deep Blending [18], and 70 images in 3 batches for Tanks&Temples [24]. We constrain the value range of $\lambda_{reg}$ for stable updates. We define it in $[1e{-}4,1e{4}]$ for Deep Blending [18] and Tanks&Temples [24] and in the interval $[1e{-}4,1e{-}2]$ for MipNeRF360 [4].
@@ 4.1 Comparison to Baselines
Taming-3DGS [27] 0.810 27.44 0.224 684 0.847 23.79 0.176 463 0.902 29.71 0.242 590
+ Ours 0.814 27.41 0.221 589 0.844 23.72 0.183 393 0.902 29.72 0.249 540
Table 1: Quantitative comparison of our method and baselines. By adding our method to baselines, we accelerate the optimization time by 30% on average while achieving the same quality. We can combine our method with others, that improve runtime along different axes. This demonstrates that our method offers an orthogonal improvement, i.e., the LM optimizer can be plugged into many existing methods.
@@
Taming-3DGS [27] after 456s Taming-3DGS + Ours after 454s Ground-Truth Images
@@
We report our main quantitative results in Tab. 1. Our LM optimizer can be added to all baseline implementations and accelerates the optimization runtime by 30% on average. The reconstructions show similar quality across all metrics and datasets, highlighting that our method arrives at similar local minima, just faster. We also provide a per-scene breakdown of these results in the supplemental material. On average our method consumes 53 GB of GPU memory on all datasets. In contrast, the baselines do not use an extra cache and only require between 6-11 GB of memory. This showcases the runtime-memory tradeoff of our approach.
@@ 4.2 Ablations
Is the L1/SSIM objective important?

### v2

@@ 18 $\mathbf{x}_{i+1}{=}\mathbf{x}_{i}{+}\alpha_{i}\mathbf{p}_{i}$
19 $\mathbf{r}_{i+1}{=}\mathbf{r}_{i}{-}\alpha_{i}\mathbf{g}_{i}$
@@ 22 $\mathbf{p}_{i+1}=\mathbf{z}_{i+1}+\beta_{i}\mathbf{p}_{i}$
Most of the work in every PCG iteration is consumed by calculating the matrix-vector product $\mathbf{g}_{i}{=}\mathbf{J}^{T}\mathbf{J}\mathbf{p}_{i}$. We compute it by first calculating $\mathbf{u}_{i}{=}\mathbf{J}\mathbf{p}_{i}$ and then $\mathbf{g}_{i}{=}\mathbf{J}^{T}\mathbf{u}_{i}$. Calculating the non-zero values of $\mathbf{J}$ requires backpropagating from the residuals through the $\alpha$-blending (Eq. 1) and splat projection steps to the Gaussian parameters. The tile-based rasterizer of 3DGS [23] performs this calculation using a per-pixel parallelization. That is, every thread handles one ray, stepping backwards along all splats that this ray hit. We found that this parallelization is too slow for an efficient PCG implementation. The reason is the repetition of the ray marching: per PCG iteration we do it once for $\mathbf{u}_{i}$ and once for $\mathbf{g}_{i}$. As a consequence, the same intermediate $\alpha$-blending states (i.e., $T_{s}$, $\frac{\partial c}{\partial\alpha_{s}}$, $\frac{\partial c}{\partial c_{s}}$ for every splat $s$ along the ray) are re-calculated multiple (up to 18) times.
Cache-driven parallelization
@@
Figure 3: Parallelization Strategy And Caching Scheme. We implement the PCG algorithm with efficient CUDA kernels, that use a gradient cache to calculate Jacobian-vector products. Left: before PCG starts, we create the gradient cache following the per-pixel parallelization of 3DGS [23]. Afterwards, we sort the cache by Gaussians to ensure coalesced read accesses. Right: the cache decouples splats along rays, which allows us to parallelize per-pixel-per-splat when computing $\mathbf{u}=\mathbf{J}\mathbf{p}$ and $\mathbf{g}=\mathbf{J}^{T}\mathbf{u}$ during PCG.
We propose to change the parallelization to per-pixel-per-splat (summarized in Fig. 3). That is, one thread handles all residuals of one ray for one splat. Each entry of $\mathbf{J}$ is the gradient from a residual $r$ (either of the L1 or SSIM terms) to a Gaussian parameter $x_{i}$. Conceptually, this can be computed in three stages:
$\frac{\partial r}{\partial x_{i}}=\frac{\partial r}{\partial c}\frac{\partial c}{\partial s}\frac{\partial s}{\partial x_{i}}$ (7)
@@
where $\frac{\partial r}{\partial c}$ denotes the gradient from the residual to the rendered color, $\frac{\partial c}{\partial s}$ from the color to the projected splat, and $\frac{\partial s}{\partial x_{i}}$ from the splat to the Gaussian parameter. The first and last factors of Eq. 7 can be computed independently for each residual and splat respectively, which allows for an efficient parallelization. Similarly, we can calculate $\frac{\partial c}{\partial s}$ independently, if we have access to $T_{s}$ and $\frac{\partial c}{\partial\alpha_{s}}$. Instead of looping over all splats along a ray multiple times, we cache these quantities once (Fig. 3 left). When calculating $\mathbf{u}_{i}$ or $\mathbf{g}_{i}$, we then read these values from the cache (Fig. 3 right). This allows us to parallelize over all splats in all pixels, which drastically accelerates the runtime. The cache size is controlled by how many images (rays) we process in each PCG iteration and how many splats contribute to the final color along each ray. We propose an efficient subsampling scheme that limits the cache size to the available budget.
3DGS uses the structural similarity index measure (SSIM) as loss term during optimization (Eq. 2). In SSIM, the local neighborhood of every pixel gets convolved with Gaussian kernels to obtain the final per-pixel score [45]. We calculate $\frac{\partial r}{\partial c}$ for the SSIM residuals by backpropagating the per-pixel scores to the center pixels (ignoring the contribution to other pixels in the local neighborhood). This allows us to keep rays independent of each other thereby allowing for an efficient parallelization. We implement it following the derivation of Zhao et al[52].
Mapping of PCG to CUDA kernels We cache all gradients $\frac{\partial c}{\partial s}$ using the buildCache operation. Following the implementation of the differentiable rasterizer in 3DGS [23], it uses the per-pixel parallelization and calculates the gradient update $\mathbf{b}{=}-\mathbf{J}^{T}\mathbf{F}$. For coalesced read and write accesses, we first store the cache sorted by pixels (Fig. 3 left). Afterwards, we re-sort it by Gaussians using the sortCacheByGaussians kernel. We use the Jacobi preconditioner $\mathbf{M}^{-1}{=}1/\text{diag}(\mathbf{J}^{T}\mathbf{J})$ and calculate it once using the per-pixel-per-splat parallelization in the diagJTJ kernel. The inner PCG loop involves two kernels that are accelerated by our novel parallelization scheme. First, applyJ computes $\mathbf{u}{=}\mathbf{J}\mathbf{p}$, which we implement as a per-pixel sum aggregation. Afterwards, applyJT computes $\mathbf{g}{=}\mathbf{J}^{T}\mathbf{u}$. This per-Gaussian sum can be efficiently aggregated using warp reductions. We compute the remaining vector-vector terms of Algorithm 1 directly in PyTorch [38]. We refer to the supplementary material for more details.
@@
Image Subsampling Scheme Our cache consumes additional GPU memory. For high resolution images in a dense reconstruction setup, the number of rays and thus the cache size can grow too large. To this end, we split the images into batches and solve the normal equations independently, following  Eq. 4. This allows us to store the cache only for one batch at a time. Concretely, for $\text{n}_{\text{b}}$ batches, we obtain $\text{n}_{\text{b}}$ update vectors and combine them in a weighted mean:
$\Delta=\sum_{i=1}^{\text{n}_{\text{b}}}\frac{\mathbf{M}_{i}\Delta_{i}}{\sum_{k=1}^{n}\mathbf{M}_{k}}$ (8)
@@
where we use the inverse of the PCG preconditioner $\mathbf{M}_{i}{=}\text{diag}(\mathbf{J}_{i}^{T}\mathbf{J}_{i})$ as the weights. We refer to the supplementary material for a derivation of the weights. These weights balance the importance of update vectors across batches based on how much each Gaussian parameter contributed to the rendered colors in the respective images. This subsampling scheme allows us to control the cache size relative to the number of images in a batch. In practice, we choose batch sizes of 25-70 images and up to $\text{n}_{\text{b}}{=}4$ batches per LM iteration. We either select the images at random or, if the scene was captured along a smooth trajectory, in a strided fashion to maximize scene coverage in all batches.
@@
Our pipeline utilizes the LM optimizer in the second stage of 3DGS optimization (see Fig. 2). Before that, we run the ADAM optimizer to obtain an initialization of the Gaussian parameters. We compare this against running our LM optimizer directly on the Gaussian initialization obtained from the SfM point cloud (following [23]). Fig. 4 shows, that our LM converges faster for better initialized Gaussians and eventually beats pure ADAM. In contrast, running it directly on the SfM initialization is slower. This demonstrates that quasi second-order solvers like ours are well-known to be more sensitive to initialization. In other words, gradient descent makes rapid progress in the beginning, but needs more time to converge to final Gaussian parameters. The additional compute overhead of our LM optimization is especially helpful to converge more quickly. This motivates us to split the method in two stages. It also allows us to complete the densification of the Gaussians before employing the LM optimizer, which simplifies the implementation.
@@ 4 Results
Figure 4: Comparison of initialization iterations. In our first stage, we initialize the Gaussians with gradient descent for K iterations, before finetuning with our LM optimizer. After $\text{K}{=}6000$ or $\text{K}{=}8000$ iterations, our method converges faster than the baseline. With less iterations, pure LM is slower, which highlights the importance of our two stage approach. Results reported on the GARDEN scene from MipNeRF360 [33] without densification.
Baselines We compare our LM optimizer against ADAM in multiple reference implementations of 3DGS. This shows, that our method is compatible with other runtime improvements. In other words, we can swap out the optimizer and retain everything else. Concretely, we compare against the original 3DGS [23], its reimplementation “gsplat” [48], and DISTWAR [12]. Additionally, we compare against Taming-3DGS [32] by utilizing their “budgeted” approach as the fastest baseline in terms of runtime. We run all baselines for 30K iterations with their default hyperparameters.
Datasets / Metrics We benchmark our runtime improvements on three established datasets: Tanks&Temples [27], Deep Blending [19], and MipNeRF360 [4]. These datasets contain in total 13 scenes that cover bounded indoor and unbounded outdoor environments. We fit all scenes for every method on the same NVIDIA A100 GPU using the train/test split as proposed in the original 3DGS [23] publication. To measure the quality of the reconstruction, we report peak signal-to-noise ratio (PSNR), structural similarity (SSIM), and perceptual similarity (LPIPS) [51] averaged over all test images. Additionally, we report the optimization runtime and the maximum amount of consumed GPU memory.
Implementation Details For our main results, we run the first stage for 20K iterations with the default hyperparameters of the respective baseline. The densification is completed after 15K iterations. Afterwards, we only have to run 5 LM iterations with 8 PCG iterations each to converge on all scenes. This showcases the efficiency of our optimizer. Since the image resolutions are different for every dataset, we select the batch-size and number of batches such that the consumed memory for caching is similar. We select 25 images in 4 batches for MipNeRF360 [4], 25 images in 3 batches for Deep Blending [19], and 70 images in 3 batches for Tanks&Temples [27]. We constrain the value range of $\lambda_{\text{reg}}$ for stable updates. We define it in $1\mathrm{e}{-4},1\mathrm{e}{4}]$ for Deep Blending [19] and Tanks&Temples [27] and in the interval $1\mathrm{e}{-4},1\mathrm{e}{-2}]$ for MipNeRF360 [4].
@@ 4.1 Comparison To Baselines
Taming-3DGS [32] 0.793 27.14 0.260 566 0.833 23.76 0.209 366 0.900 29.84 0.274 447
+ Ours 0.791 27.13 0.260 453 0.832 23.72 0.209 310 0.901 29.91 0.275 347
Table 1: Quantitative comparison of our method and baselines. By adding our method to baselines, we accelerate the optimization time by 20% on average while achieving the same quality. We can combine our method with others, that improve runtime along different axes. This demonstrates that our method offers an orthogonal improvement, i.e., the LM optimizer can be plugged into many existing methods.
@@
Taming-3DGS [32] after 328s Taming-3DGS + Ours after 324s Ground-Truth Images
@@
We report our main quantitative results in Tab. 1. Our LM optimizer can be added to all baseline implementations and accelerates the optimization runtime by 20% on average. The reconstructions show similar quality across all metrics and datasets, highlighting that our method arrives at similar local minima, just faster. We also provide a per-scene breakdown of these results in the supplementary material. On average our method consumes 53 GB of GPU memory on all datasets. In contrast, the baselines do not use an extra cache and only require between 6-11 GB of memory. This showcases the runtime-memory tradeoff of our approach.
@@ 4.2 Ablations
Is the L1/SSIM objective important? We utilize the same objective in our LM optimizer as in the original 3DGS implementation, namely the L1 and SSIM loss terms (Eq. 2). Since LM energy terms are defined as a sum of squares, we adopt the square root formulation of these loss terms to arrive at an identical objective (Eq. 3). We compare this choice against fitting the Gaussians with only an L2 loss, that does not require taking a square root. Concretely, we compare the achieved quality and runtime of LM against ADAM for both the L2 loss and the L1 and SSIM losses. As can be seen in Tab. 2, we achieve faster convergence and similar quality in both cases. However, the achieved quality is inferior for both LM and ADAM when only using the L2 loss. This highlights the importance of the L1 and SSIM loss terms and why we adopt them in our method as well. We show in the supplementary material, that computing these loss terms instead of the simpler L2 residuals does not negatively impact the efficiency of our CUDA kernels.
How many images per batch are necessary? The key hyperparameters in our model are the number of images in a batch and how many batches to choose for every LM iteration (Sec. 3.3). This controls the runtime of one iteration and how much GPU memory our optimizer consumes. We compare different numbers of images in Tab. 3 on the NeRF-Synthetic [33] dataset in a single batch per LM iteration, i.e., $\text{n}_{\text{b}}{=}1$. Using the full dataset (100 images) produces the best results. Decreasing the number of images in a batch results in only slightly worse quality, but also yields faster convergence and reduces GPU memory consumption linearly down to 15GB for 40 images. This demonstrates that subsampling images does not negatively impact the convergence of the LM optimizer in our task.
Are we better than multi-view ADAM? Our method converges with fewer iterations than baselines. Concretely, we require only 5-10 additional LM iterations after the initialization, whereas ADAM runs for another 10K iterations. We increase the batch-size (number of images) for the baselines, such that the same number of multi-view constraints are observed for the respective update steps. However, as can be seen in Tab. 4, the achieved quality is worse for ADAM after the same number of iterations. When running for more iterations, ADAM eventually converges to similar quality, but needs more time. This highlights the efficiency of our optimizer: since we solve the normal equations in Eq. 3, one LM iteration makes a higher quality update step than ADAM which only uses the gradient direction.
4.3 Runtime Analysis
We analyze the runtime of our LM optimizer across multiple iterations in Fig. 6. The runtime is dominated by solving Eq. 4 with PCG and building the cache (Sec. 3.3). Sorting the cache, rendering the selected images, and the line search (Eq. 5) are comparatively faster. During PCG, we run the applyJ and applyJT kernels up to 9 times, parallelizing per-pixel-per-splat. In contrast, we run the buildCache kernel once, parallelizing per-pixel, which is only marginally faster. This shows the advantage of our proposed parallelization scheme: the same Jacobian-vector product runs much faster. We also provide a detailed profiling analysis of our kernels in the supplementary material.
4.4 Limitations